Use the balanced reaction below to answer the following question.

Al4C3(s) + 12 H2O(l) → 4 Al(OH)3(s) + 3 CH4(g)

Given 6.99 grams of Al4C3 and 7.78 grams of H2O, calculate the number of grams of Al(OH)3 that are produced.
Answer #1

All 3 (5) + 12 H2010) –> 4 A1 (0H), (3) + 3 CHA (9) Molao mass of Alufs (4x 27)+ (3 xia) = 108 + 3b = 1444 Molaro mass of H2O12 H2O 216 As per Aluc geachion, lių given a 7-18 no 144x7.78 -5.1869 Only 5.186g of Alu will The total Alloh), produced = reThr answer is 11.237g